# Improving LLM First-Token Predictions in Multiple-Choice Question Answering via Prefilling Attack

### v1

@@ 4.6 Symbolic Output Structure Analysis
While standard accuracy metrics quantify how often a model fails, they offer limited insight into how models fail. To deepen our understanding of model behavior, we analyze auxiliary statistics that capture qualitative patterns in error generation.
Specifically, to analyze robustness against first-token misinterpretation, we introduce a new metric that answers the following question: When the model predicts a valid symbolic answer (e.g., ‘A’, ‘B’), how consistent is the continuation? Ideally, an aligned model should either stop after the symbolic token or follow it with a predictable structure. In contrast, inconsistent or verbose continuations may indicate that the answer token was generated as part of a grammatical phrase rather than as an intended choice. We capture this behavior using the Continuation Diversity (CD), defined as the number of distinct second tokens that follow a valid first-token prediction, in relation to the calculated First-Token Validity Rate (FTVR). Formally, it is computed as $\text{CD}=\frac{{S}}{\text{FTVR}},$ where ${S}$ is the number of unique second tokens that follow a valid first-token prediction. Low CD suggests that the model reliably adheres to a symbolic format (e.g., consistently producing ‘A.’), whereas high CD may signal poor format consistency and potential misinterpretation.
@@
Table 5: Evaluation of symbolic output structure across four MCQA benchmarks in terms of CD and FTVR. Prefilling consistently increases FTVR and lowers CD, indicating that it reduces misinterpretation errors and promotes more stable symbolic completions.
5 Conclusions
This work demonstrates the significant benefits of using a prefilling attack as a simple yet powerful intervention to enhance the reliability of FTP evaluation in multiple-choice question-answering tasks. By adding a structured prefix to the model response template, we effectively mitigate issues of first-token misalignment and misinterpretation, steering the model toward more accurate and consistent predictions. Experiments across several MCQA benchmarks show that this method not only improves alignment between the predicted token and the correct answer but also boosts overall accuracy and calibration. Moreover, a key finding is that prefilling leads to FTP accuracies that are comparable to those obtained through open-ended generation evaluated via GPT or Llama classifiers. Notably, these gains are achieved without fine-tuning or any model modifications. This approach presents a promising path for enhancing LLM performance in choice-based evaluation scenarios.
Limitations
Acknowledgments
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources. This work has been supported by the projects MINERVA (GA No 101182737), ELLIOT (GA No 101214398), both funded by the European Union.
Appendix A Appendix
A.1 Analysis Across Multiple Templates
In addition to the main experiments, which utilize the default prefilling template (i.e., ‘Given the question and the possible options, my answer is: ’), we conduct a supplementary analysis to assess the robustness of models performance across template variations. Specifically, we evaluate the models using a set of 10 distinct templates, each phrased differently but conveying the same task. For this analysis, we report the average accuracy and standard deviation across the 10 templates to capture both overall performance and sensitivity to prompt formulation. The set of templates used in this analysis alternates short, concise wording to longer, more context-aware versions, like:
@@ A.1 Analysis Across Multiple Templates
  • Given the context, phrasing, and underlying assumptions in both the question and its options, I determine the most fitting response to be:
The average accuracies, together with their standard deviations, are reported in Table 6, alongside the results already displayed in Table 1. From the results, we can observe that these average accuracies are generally comparable to those obtained using the single default template. These results demonstrate that the prefilling strategy is robust across different template formulations and is not highly sensitive to template phrasing. In a few rare cases, however, significant improvements are observed. For example, Gemma-7B achieves an accuracy 5.9 percentage points higher than the default template on the Moral Stories benchmark, while Phi-4-14B shows a 5.8-point gain on the LogiQA benchmark. These results suggest that the default template may not always be the most effective choice.
This variability aligns with the understanding that no single template style is optimal across all models. Differences in architecture, pretraining corpora, and alignment strategies influence how models interpret and get affected by certain tokens. As a result, a template that benefits one model may be suboptimal for another.
Moreover, we find that longer templates can exert a stronger influence on the model generation. While this added guidance is generally beneficial, leading to better task alignment, it may also introduce unintended biases or dilute the task signal in others. These variations open opportunities for users to tailor template styles to the specific model they are using, while retaining confidence that reasonable rephrasings do not significantly degrade performance.
@@
MMLU RACE OBQA SIQA MS CQA SC HS MC-T SciQ MQA ARC-E LQA
Llama-3.1-8B 63.1 78.1 71.2 70.9 87.0 72.7 93.2 52.2 82.1 97.1 24.0 90.2 28.9
+ prefilling (default) 68.4 83.2 84.8 72.3 93.2 76.5 96.4 69.0 82.5 98.3 33.8 94.7 34.0
+ prefilling (avg) 66.9±$\pm 8.6\mathrm{e}^{-3}$ 82.9±$\pm 5.9\mathrm{e}^{-3}$ 81.7±$\pm 1.4\mathrm{e}^{-2}$ 71.8±$\pm 6.6\mathrm{e}^{-3}$ 91.3±$\pm 2.8\mathrm{e}^{-2}$ 76.6±$\pm 5.6\mathrm{e}^{-3}$ 96.4±$\pm 7.7\mathrm{e}^{-3}$ 66.0±$\pm 2.8\mathrm{e}^{-2}$ 79.0±$\pm 5.5\mathrm{e}^{-2}$ 98.1±$\pm 1.4\mathrm{e}^{-3}$ 34.9±$\pm 9.6\mathrm{e}^{-3}$ 92.7±$\pm 3.9\mathrm{e}^{-3}$ 33.3±$\pm 6.2\mathrm{e}^{-3}$
Qwen-2-7B 68.9 87.0 83.8 75.6 88.1 79.4 97.1 53.4 88.3 97.5 29.8 94.3 33.0
+ prefilling (default) 69.1 86.7 84.8 76.1 92.6 79.4 97.3 65.2 86.1 97.4 36.5 94.7 35.5
+ prefilling (avg) 68.9±$\pm 8.6\mathrm{e}^{-3}$ 86.6±$\pm 3.4\mathrm{e}^{-3}$ 84.4±$\pm 1.0\mathrm{e}^{-2}$ 76.1±$\pm 6.0\mathrm{e}^{-3}$ 91.7±$\pm 4.8\mathrm{e}^{-2}$ 79.5±$\pm 6.1\mathrm{e}^{-3}$ 97.5±$\pm 2.6\mathrm{e}^{-3}$ 65.3±$\pm 6.0\mathrm{e}^{-2}$ 86.7±$\pm 4.5\mathrm{e}^{-2}$ 97.5±$\pm 3.7\mathrm{e}^{-3}$ 37.3±$\pm 1.4\mathrm{e}^{-2}$ 94.7±$\pm 1.9\mathrm{e}^{-3}$ 33.1±$\pm 6.2\mathrm{e}^{-3}$
Gemma-7B 45.9 50.4 40.6 34.2 77.5 51.6 69.6 36.3 60.1 89.0 24.1 55.9 26.4
+ prefilling (default) 52.4 70.5 71.2 65.2 64.3 68.6 92.0 56.3 91.0 95.1 25.1 86.7 26.1
+ prefilling (avg) 52.7±$\pm 3.8\mathrm{e}^{-3}$ 70.3±$\pm 2.5\mathrm{e}^{-3}$ 70.9±$\pm 7.2\mathrm{e}^{-3}$ 65.3±$\pm 4.5\mathrm{e}^{-3}$ 70.4±$\pm 7.1\mathrm{e}^{-2}$ 68.3±$\pm 4.7\mathrm{e}^{-3}$ 92.1±$\pm 6.5\mathrm{e}^{-3}$ 55.5±$\pm 2.4\mathrm{e}^{-2}$ 89.6±$\pm 2.3\mathrm{e}^{-2}$ 94.9±$\pm 2.6\mathrm{e}^{-3}$ 25.7±$\pm 1.2\mathrm{e}^{-2}$ 86.8±$\pm 3.5\mathrm{e}^{-3}$ 26.2±$\pm 6.5\mathrm{e}^{-3}$
Gemma-2-9B 33.9 49.5 38.8 56.3 85.9 39.6 97.6 43.0 72.2 31.8 27.8 38.6 31.0
+ prefilling (default) 72.1 86.3 89.8 74.2 87.3 79.0 97.9 69.8 78.5 98.3 34.0 96.6 35.8
+ prefilling (avg) 71.6±$\pm 3.5\mathrm{e}^{-3}$ 86.4±$\pm 3.3\mathrm{e}^{-3}$ 89.4±$\pm 1.1\mathrm{e}^{-2}$ 74.4±$\pm 2.1\mathrm{e}^{-3}$ 88.0±$\pm 4.1\mathrm{e}^{-2}$ 78.9±$\pm 4.7\mathrm{e}^{-3}$ 97.8±$\pm 1.9\mathrm{e}^{-3}$ 70.9±$\pm 5.2\mathrm{e}^{-2}$ 77.3±$\pm 2.1\mathrm{e}^{-2}$ 98.3±$\pm 1.1\mathrm{e}^{-3}$ 34.5±$\pm 2.0\mathrm{e}^{-2}$ 96.6±$\pm 1.0\mathrm{e}^{-3}$ 34.1±$\pm 1.1\mathrm{e}^{-2}$
Zephyr-7B 57.5 86.5 73.0 67.5 83.6 66.0 96.0 35.0 69.1 88.5 23.3 95.6 32.1
+ prefilling (default) 58.8 87.3 70.8 68.5 86.2 69.0 96.0 36.1 93.3 93.1 24.3 98.1 40.6
+ prefilling (avg) 58.1±$\pm 2.1\mathrm{e}^{-3}$ 73.4±$\pm 2.6\mathrm{e}^{-3}$ 71.6±$\pm 1.1\mathrm{e}^{-2}$ 67.5±$\pm 6.8\mathrm{e}^{-3}$ 85.7±$\pm 3.9\mathrm{e}^{-2}$ 68.5±$\pm 1.4\mathrm{e}^{-2}$ 96.3±$\pm 3.9\mathrm{e}^{-3}$ 36.3±$\pm 2.0\mathrm{e}^{-2}$ 91.0±$\pm 4.2\mathrm{e}^{-2}$ 93.8±$\pm 6.5\mathrm{e}^{-3}$ 24.7±$\pm 1.7\mathrm{e}^{-2}$ 87.8±$\pm 3.5\mathrm{e}^{-3}$ 33.1±$\pm 1.3\mathrm{e}^{-2}$
Ministral-8B 62.1 84.7 82.4 74.7 87.3 74.3 97.4 81.1 91.5 97.4 23.1 93.6 28.4
+ prefilling (default) 63.9 84.7 85.2 75.9 89.9 75.0 98.2 86.5 91.9 98.0 24.5 93.7 34.6
+ prefilling (avg) 63.5±$\pm 4.2\mathrm{e}^{-3}$ 84.3±$\pm 6.5\mathrm{e}^{-3}$ 84.2±$\pm 1.2\mathrm{e}^{-2}$ 75.5±$\pm 5.4\mathrm{e}^{-3}$ 90.0±$\pm 3.1\mathrm{e}^{-2}$ 75.3±$\pm 4.4\mathrm{e}^{-3}$ 98.0±$\pm 1.6\mathrm{e}^{-3}$ 85.9±$\pm 1.0\mathrm{e}^{-2}$ 91.6±$\pm 2.0\mathrm{e}^{-2}$ 97.6±$\pm 2.6\mathrm{e}^{-3}$ 24.5±$\pm 1.4\mathrm{e}^{-2}$ 93.4±$\pm 2.7\mathrm{e}^{-3}$ 32.2±$\pm 1.1\mathrm{e}^{-2}$
Mistral-Nemo-12B 65.1 82.7 80.6 74.2 78.0 74.1 97.3 51.6 92.4 96.8 23.5 92.6 28.7
+ prefilling (default) 66.0 83.1 80.8 75.8 80.9 75.4 96.9 76.4 93.3 97.2 25.4 93.1 32.0
+ prefilling (avg) 65.0±$\pm 7.4\mathrm{e}^{-3}$ 83.0±$\pm 1.0\mathrm{e}^{-2}$ 79.8±$\pm 1.2\mathrm{e}^{-2}$ 75.1±$\pm 4.6\mathrm{e}^{-3}$ 82.4±$\pm 9.0\mathrm{e}^{-2}$ 74.3±$\pm 1.1\mathrm{e}^{-2}$ 97.0±$\pm 3.4\mathrm{e}^{-3}$ 73.1±$\pm 3.4\mathrm{e}^{-2}$ 91.7±$\pm 3.2\mathrm{e}^{-2}$ 97.2±$\pm 4.0\mathrm{e}^{-3}$ 25.7±$\pm 1.5\mathrm{e}^{-2}$ 92.5±$\pm 7.9\mathrm{e}^{-3}$ 31.4±$\pm 6.9\mathrm{e}^{-3}$
Phi-4-14B 76.4 73.3 84.0 71.7 83.4 72.5 98.2 61.2 82.5 95.4 25.0 87.8 31.9
+ prefilling (default) 79.7 73.4 90.0 77.3 93.2 79.6 98.7 87.8 86.1 98.3 45.0 87.8 34.6
+ prefilling (avg) 79.0±$\pm 2.1\mathrm{e}^{-3}$ 87.2±$\pm 2.7\mathrm{e}^{-3}$ 89.6±$\pm 1.4\mathrm{e}^{-2}$ 76.5±$\pm 1.1\mathrm{e}^{-2}$ 94.1±$\pm 3.7\mathrm{e}^{-2}$ 79.4±$\pm 6.2\mathrm{e}^{-3}$ 98.4±$\pm 3.5\mathrm{e}^{-3}$ 86.7±$\pm 1.7\mathrm{e}^{-2}$ 81.2±$\pm 1.0\mathrm{e}^{-1}$ 98.0±$\pm 4.4\mathrm{e}^{-3}$ 45.2±$\pm 2.8\mathrm{e}^{-2}$ 98.1±$\pm 2.3\mathrm{e}^{-3}$ 40.4±$\pm 2.1\mathrm{e}^{-2}$
A.2 Open-Ended Generation

### v2

@@ 4.6 Symbolic Output Structure Analysis
While standard accuracy metrics quantify how often a model fails, they offer limited insight into how models fail. To deepen our understanding of model behavior, we analyze auxiliary statistics that capture qualitative patterns in error generation. Specifically, to analyze robustness against first-token misinterpretation, we introduce a new metric that answers the following question: When the model predicts a valid symbolic answer (e.g., ‘A’, ‘B’), how consistent is the continuation? Ideally, an aligned model should either stop after the symbolic token or follow it with a predictable structure. In contrast, inconsistent or verbose
Table 4: Evaluation of symbolic output structure across four MCQA benchmarks in terms of CD and FTVR. Prefilling consistently increases FTVR and lowers CD, indicating fewer misinterpretation errors and more stable symbolic completions.
@@
continuations may indicate that the answer token was generated as part of a grammatical phrase rather than as an intended choice. We capture this behavior using the Continuation Diversity (CD), defined as the number of distinct second tokens that follow a valid first-token prediction, in relation to the calculated First-Token Validity Rate (FTVR). Formally, it is computed as $\text{CD}=\frac{{S}}{\text{FTVR}},$ where S{S} is the number of unique second tokens that follow a valid first-token prediction. Low CD suggests that the model reliably adheres to a symbolic format (e.g., consistently producing ‘A.’), whereas high CD may signal poor format consistency and potential misinterpretation.
5 Conclusion
This work demonstrates the significant benefits of using output prefilling to enhance the reliability of FTP evaluation in MCQA tasks. By adding a structured prefix to the model response template, we substantially improve issues of first-token misalignment and misinterpretation, steering the model toward more accurate and consistent predictions. Experiments across several MCQA benchmarks show that this method not only improves alignment between the predicted token and the correct answer but also boosts overall accuracy and calibration, with FTP accuracies comparable to those obtained via open-ended generation evaluated by external GPT, Llama, or xFinder classifiers. Notably, these gains are achieved without fine-tuning or any model modifications. To our knowledge, this is the first work to rigorously quantify the effectiveness of prefilling with modern general-purpose LLMs, highlighting its promise as a lightweight strategy for choice-based evaluation scenarios.
5.0.1 Acknowledgments
This work has been supported by the EU Horizon projects ELIAS (No. 101120237) and ELLIOT (No. 101214398), and by the EuroHPC JU project MINERVA (No. 101182737). We also acknowledge CINECA for the availability of high-performance computing resources under the ISCRA initiative.
[…]
Samuele Poppi3This work was carried out primarily while the author was affiliated with the University of Modena and Reggio Emilia, Italy.Zheng Xin Yong4Diego Garcia-Olano5
Marcella Cornia1Lorenzo Baraldi1Rita Cucchiara1
Appendix 0.A Analysis Across Multiple Templates
In addition to the main experiments, which utilize the default prefilling template (i.e., ‘Given the question and the possible options, my answer is: ’), we conduct an additional analysis to assess the robustness of models performance across template variations. Specifically, we evaluate the models using a set of 10 distinct templates, each phrased differently but conveying the same task. For this analysis, we report the average accuracy and standard deviation across the 10 templates to capture both overall performance and sensitivity to prompt formulation.
The set of templates used in this analysis alternates short, concise wording to longer, more context-aware versions, like:
@@ Appendix 0.A Analysis Across Multiple Templates
  • Given the context and underlying assumptions in both the question and its options, I determine the most fitting response to be:
The average accuracies, together with their standard deviations, are reported in Table 5, alongside the results already displayed in Table 1 of the main paper. From the results, we can observe that these average accuracies are generally comparable to those obtained using the single default template. These results demonstrate that the prefilling strategy is robust across different template formulations and is not highly sensitive to template phrasing. In a few rare cases, however, significant improvements are observed. For example, Gemma-7B achieves an accuracy 5.9 percentage points higher than the default template on the Moral Stories benchmark, while Phi-4-14B shows a 5.8-point gain on the LogiQA benchmark. These results suggest that the default template may not always be the most effective choice.
This variability aligns with the understanding that no single template style is optimal across all models. Differences in architecture, pretraining corpora, and alignment strategies influence how models interpret and get affected by certain tokens. As a result, a template that benefits one model may be suboptimal for another. Moreover, we find that longer templates can exert a stronger influence on the model generation. While this added guidance is generally beneficial, leading to better task alignment, it may also introduce unintended biases or dilute the task signal in others. These variations open opportunities for users to tailor template styles to the specific model they are using, while retaining confidence that reasonable rephrasings do not significantly degrade performance.
@@
MMLU RACE OBQA SIQA MS CQA SC HS MC-T SciQ MQA ARC-E ARC-C LQA
Llama-3.1-8B 63.1 78.1 71.2 70.9 87.0 72.7 93.2 52.2 82.1 97.1 24.0 90.2 52.7 28.9
+ prefilling (default) 68.4 83.2 84.8 72.3 93.2 76.5 96.4 69.0 82.5 98.3 33.8 94.7 60.2 34.0
+ prefilling (avg) 66.9 82.9 81.7 71.8 91.3 76.6 96.4 66.0 79.0 98.1 34.9 92.7 58.8 33.3
$\pm 8.6\mathrm{e}^{-3}$ $\pm 5.9\mathrm{e}^{-3}$ $\pm 1.4\mathrm{e}^{-2}$ $\pm 6.6\mathrm{e}^{-3}$ $\pm 2.8\mathrm{e}^{-2}$ $\pm 5.6\mathrm{e}^{-3}$ $\pm 7.7\mathrm{e}^{-3}$ $\pm 2.8\mathrm{e}^{-2}$ $\pm 5.5\mathrm{e}^{-2}$ $\pm 1.4\mathrm{e}^{-3}$ $\pm 9.6\mathrm{e}^{-3}$ $\pm 3.9\mathrm{e}^{-3}$ $\pm 9.0\mathrm{e}^{-3}$ $\pm 6.2\mathrm{e}^{-3}$
Qwen-2-7B 68.9 87.0 83.8 75.6 88.1 79.4 97.1 53.4 88.3 97.5 29.8 94.3 48.0 33.0
+ prefilling (default) 69.1 86.7 84.8 76.1 92.6 79.4 97.3 65.2 86.1 97.4 36.5 94.7 56.6 35.5
+ prefilling (avg) 68.9 86.6 84.4 76.1 91.7 79.5 97.5 65.3 86.7 97.5 37.3 94.7 55.9 33.1
$\pm 8.6\mathrm{e}^{-3}$ $\pm 3.4\mathrm{e}^{-3}$ $\pm 1.0\mathrm{e}^{-2}$ $\pm 6.0\mathrm{e}^{-3}$ $\pm 4.8\mathrm{e}^{-2}$ $\pm 6.1\mathrm{e}^{-3}$ $\pm 2.6\mathrm{e}^{-3}$ $\pm 6.0\mathrm{e}^{-2}$ $\pm 4.5\mathrm{e}^{-2}$ $\pm 3.7\mathrm{e}^{-3}$ $\pm 1.4\mathrm{e}^{-2}$ $\pm 1.9\mathrm{e}^{-3}$ $\pm 2.0\mathrm{e}^{-2}$ $\pm 6.2\mathrm{e}^{-3}$
Gemma-7B 45.9 50.4 40.6 34.2 77.5 51.6 69.6 36.3 60.1 89.0 24.1 55.9 43.9 26.4
+ prefilling (default) 52.4 70.5 71.2 65.2 64.3 68.6 92.0 56.3 91.0 95.1 25.1 86.7 49.1 26.1
+ prefilling (avg) 52.7 70.3 70.9 65.3 70.4 68.3 92.1 55.5 89.6 94.9 25.7 86.8 49.6 26.2
$\pm 3.8\mathrm{e}^{-3}$ $\pm 2.5\mathrm{e}^{-3}$ $\pm 7.2\mathrm{e}^{-3}$ $\pm 4.5\mathrm{e}^{-3}$ $\pm 7.1\mathrm{e}^{-2}$ $\pm 4.7\mathrm{e}^{-3}$ $\pm 6.5\mathrm{e}^{-3}$ $\pm 2.4\mathrm{e}^{-2}$ $\pm 2.3\mathrm{e}^{-2}$ $\pm 2.6\mathrm{e}^{-3}$ $\pm 1.2\mathrm{e}^{-2}$ $\pm 3.5\mathrm{e}^{-3}$ $\pm 1.5\mathrm{e}^{-2}$ $\pm 6.5\mathrm{e}^{-3}$
Gemma-2-9B 33.9 49.5 38.8 56.3 85.9 39.6 97.6 43.0 72.2 31.8 27.8 38.6 52.4 31.0
+ prefilling (default) 72.1 86.3 89.8 74.2 87.3 79.0 97.9 69.8 78.5 98.3 34.0 96.6 60.8 35.8
+ prefilling (avg) 71.6 86.4 89.4 74.4 88.0 78.9 97.8 70.9 77.3 98.3 34.5 96.6 59.8 34.1
$\pm 3.5\mathrm{e}^{-3}$ $\pm 3.3\mathrm{e}^{-3}$ $\pm 1.1\mathrm{e}^{-2}$ $\pm 2.1\mathrm{e}^{-3}$ $\pm 4.1\mathrm{e}^{-2}$ $\pm 4.7\mathrm{e}^{-3}$ $\pm 1.9\mathrm{e}^{-3}$ $\pm 5.2\mathrm{e}^{-2}$ $\pm 2.1\mathrm{e}^{-2}$ $\pm 1.1\mathrm{e}^{-3}$ $\pm 2.0\mathrm{e}^{-2}$ $\pm 1.0\mathrm{e}^{-3}$ $\pm 2.0\mathrm{e}^{-2}$ $\pm 1.1\mathrm{e}^{-2}$
Zephyr-7B 57.5 86.5 73.0 67.5 83.6 66.0 96.0 35.0 69.1 88.5 23.3 95.6 51.0 32.1
+ prefilling (default) 58.8 87.3 70.8 68.5 86.2 69.0 96.0 36.1 93.3 93.1 24.3 98.1 51.0 40.6
+ prefilling (avg) 58.1 73.4 71.6 67.5 85.7 68.5 96.3 36.3 91.0 93.8 24.7 87.8 52.2 33.1
$\pm 2.1\mathrm{e}^{-3}$ $\pm 2.6\mathrm{e}^{-3}$ $\pm 1.1\mathrm{e}^{-2}$ $\pm 6.8\mathrm{e}^{-3}$ $\pm 3.9\mathrm{e}^{-2}$ $\pm 1.4\mathrm{e}^{-2}$ $\pm 3.9\mathrm{e}^{-3}$ $\pm 2.0\mathrm{e}^{-2}$ $\pm 4.2\mathrm{e}^{-2}$ $\pm 6.5\mathrm{e}^{-3}$ $\pm 1.7\mathrm{e}^{-2}$ $\pm 3.5\mathrm{e}^{-3}$ $\pm 1.4\mathrm{e}^{-2}$ $\pm 1.3\mathrm{e}^{-2}$
Ministral-8B 62.1 84.7 82.4 74.7 87.3 74.3 97.4 81.1 91.5 97.4 23.1 93.6 48.1 28.4
+ prefilling (default) 63.9 84.7 85.2 75.9 89.9 75.0 98.2 86.5 91.9 98.0 24.5 93.7 49.1 34.6
+ prefilling (avg) 63.5 84.3 84.2 75.5 90.0 75.3 98.0 85.9 91.6 97.6 24.5 93.4 51.3 32.2
$\pm 4.2\mathrm{e}^{-3}$ $\pm 6.5\mathrm{e}^{-3}$ $\pm 1.2\mathrm{e}^{-2}$ $\pm 5.4\mathrm{e}^{-3}$ $\pm 3.1\mathrm{e}^{-2}$ $\pm 4.4\mathrm{e}^{-3}$ $\pm 1.6\mathrm{e}^{-3}$ $\pm 1.0\mathrm{e}^{-2}$ $\pm 2.0\mathrm{e}^{-2}$ $\pm 2.6\mathrm{e}^{-3}$ $\pm 1.4\mathrm{e}^{-2}$ $\pm 2.7\mathrm{e}^{-3}$ $\pm 3.3\mathrm{e}^{-2}$ $\pm 1.1\mathrm{e}^{-2}$
Mistral-Nemo-12B 65.1 82.7 80.6 74.2 78.0 74.1 97.3 51.6 92.4 96.8 23.5 92.6 51.3 28.7
+ prefilling (default) 66.0 83.1 80.8 75.8 80.9 75.4 96.9 76.4 93.3 97.2 25.4 93.1 54.7 32.0
+ prefilling (avg) 65.0 83.0 79.8 75.1 82.4 74.3 97.0 73.1 91.7 97.2 25.7 92.5 54.4 31.4
$\pm 7.4\mathrm{e}^{-3}$ $\pm 1.0\mathrm{e}^{-2}$ $\pm 1.2\mathrm{e}^{-2}$ $\pm 4.6\mathrm{e}^{-3}$ $\pm 9.0\mathrm{e}^{-2}$ $\pm 1.1\mathrm{e}^{-2}$ $\pm 3.4\mathrm{e}^{-3}$ $\pm 3.4\mathrm{e}^{-2}$ $\pm 3.2\mathrm{e}^{-2}$ $\pm 4.0\mathrm{e}^{-3}$ $\pm 1.5\mathrm{e}^{-2}$ $\pm 7.9\mathrm{e}^{-3}$ $\pm 2.1\mathrm{e}^{-2}$ $\pm 6.9\mathrm{e}^{-3}$
Phi-4-14B 76.4 73.3 84.0 71.7 83.4 72.5 98.2 61.2 82.5 95.4 25.0 87.8 46.4 31.9
+ prefilling (default) 79.7 73.4 90.0 77.3 93.2 79.6 98.7 87.8 86.1 98.3 45.0 87.8 60.2 34.6
+ prefilling (avg) 79.0 87.2 89.6 76.5 94.1 79.4 98.4 86.7 81.2 98.0 45.2 98.1 59.6 40.4
$\pm 2.1\mathrm{e}^{-3}$ $\pm 2.7\mathrm{e}^{-3}$ $\pm 1.4\mathrm{e}^{-2}$ $\pm 1.1\mathrm{e}^{-2}$ $\pm 3.7\mathrm{e}^{-2}$ $\pm 6.2\mathrm{e}^{-3}$ $\pm 3.5\mathrm{e}^{-3}$ $\pm 1.7\mathrm{e}^{-2}$ $\pm 1.0\mathrm{e}^{-1}$ $\pm 4.4\mathrm{e}^{-3}$ $\pm 2.8\mathrm{e}^{-2}$ $\pm 2.3\mathrm{e}^{-3}$ $\pm 3.3\mathrm{e}^{-2}$ $\pm 2.1\mathrm{e}^{-2}$
Appendix 0.B Open-Ended Generation